### v1

# A Vision-Language Pre-training Model-Guided Approach for Mitigating Backdoor Attacks in Federated Learning

Abstract
Existing backdoor defense methods in Federated Learning (FL) rely on the assumption of homogeneous client data distributions or the availability of a clean serve dataset, which limits the practicality and effectiveness. Defending against backdoor attacks under heterogeneous client data distributions while preserving model performance remains a significant challenge. In this paper, we propose a FL backdoor defense framework named CLIP-Fed, which leverages the zero-shot learning capabilities of vision-language pre-training models. By integrating both pre-aggregation and post-aggregation defense strategies, CLIP-Fed overcomes the limitations of Non-IID imposed on defense effectiveness. To address privacy concerns and enhance the coverage of the dataset against diverse triggers, we construct and augment the server dataset using the multimodal large language model and frequency analysis without any client samples. To address class prototype deviations caused by backdoor samples and eliminate the correlation between trigger patterns and target labels, CLIP-Fed aligns the knowledge of the global model and CLIP on the augmented dataset using prototype contrastive loss and Kullback-Leibler divergence. Extensive experiments on representative datasets validate the effectiveness of CLIP-Fed. Compared to state-of-the-art methods, CLIP-Fed achieves an average reduction in ASR, i.e., 2.03% on CIFAR-10 and 1.35% on CIFAR-10-LT, while improving average MA by 7.92% and 0.48%, respectively.
Codehttps://anonymous.4open.science/r/CLIP-Fed
Introduction
Figure 1: An overall framework of CLIP-Fed. We utilize the prior knowledge and cross-modal visual-language representation capabilities of CLIP to purify backdoor.
The robustness of Federated Learning (FL) (Han et al. 2023; Fang, Ye, and Du 2025) has become a growing concern, particularly under adversarial settings. The decentralized training process makes FL vulnerable to backdoor attacks (Zhuang et al. 2024), as malicious clients can poison local training data to induce the global model and associate specific trigger patterns (e.g., pixel patches or watermarks) with target labels. Due to FL’s distributed nature, the backdoor attacks become more stealthy and cause significantly more severe damage than in centralized settings (Liao et al. 2025; Lyu et al. 2025).
Existing defense strategies can be classified into three main categories. Specifically, the first type distinguishes benign and malicious clients by clustering model parameters (Nguyen et al. 2022; Mu et al. 2024; Fereidooni et al. 2024), while highly relies on the homogeneous data distribution of FL. The second depends on modified aggregation strategies such as weighted aggregation, median-based aggregation, and rank-based aggregation to achieve robust FL (Han et al. 2023; Cao et al. 2021; Yin et al. 2018). However, malicious clients can adjust attack strategies to circumvent these rules, such as the screening of poisoning parameters and hierarchical poisoning mechanisms (Zhuang et al. 2024). The third kind of methods use differential privacy to reduce the impact of backdoor attacks, but task performance is generally negatively affected (Nguyen et al. 2022; Bagdasaryan et al. 2020). Thus, most existing methods are built on overly rigid assumptions of accessing to a clean labeled validation set or homogeneous data distribution across clients, which limit their real-world utility. Moreover, existing defense methods overlook the issue of class prototype deviations caused by the presence of backdoor samples, which leads to a degradation in model performance during the backdoor defense process.
To address above challenges, we propose to leverage the rich prior knowledge and cross-modal visual-language representation capabilities of Vision-Language Pretraining models (VLPs) to defend the global model against backdoor attacks, as shown in Figure 1. CLIP (Radford et al. 2021) is a representative VLP that uses contrastive learning, and its powerful zero-shot learning capability makes it a potential tool for purifying backdoor patterns in the global model. First, we integrate both pre- and post-aggregation defense strategies to eliminate reliance on homogeneous client data distributions. Specifically, pre-aggregation model clustering filters out malicious updates, while post-aggregation purification uses a global model adjustment to enhance robustness under Non-IID conditions. To further address privacy concerns, we construct a server dataset by combining Multimodal Large Language Models (MLLMs) and frequency analysis without any client samples, which balances privacy protection, coverage of diverse backdoor triggers, and the impact of image augmentation on model utility. Finally, by using the enhanced server dataset under CLIP’s semantic guidance via prototype contrastive learning, we rectify the global model’s feature extractor to correct class prototype deviations caused by backdoor samples, while decoupling the strong correlations between trigger patterns and target labels. To this end, we align the logits distributions of CLIP and the global model through knowledge transfer to further remove residual backdoor associations.
The main contributions are summarized as follows. (1) We propose ̵CLIP-Fed, the first ̵VLP-guided FL backdoor defense framework, which uses the cross-modal representation capabilities of VLPs to mitigate backdoor attacks. CLIP-Fed combines pre-aggregation model clustering and post-aggregation backdoor purification to jointly enhance defense robustness, reducing dependence on homogeneous client data distributions. (2) To correct ̵class prototype deviations induced by backdoor samples, we propose a ̵semantically supervised global model rectification method using CLIP, which employs a ̵prototype contrastive loss and ̵Kullback-Leibler (KL) divergence to align the global model features with CLIP’s embeddings in an enhanced dataset, thereby ̵eliminating correlations between backdoor triggers and target labels. (3) Extensive evaluations on ̵three datasets and ̵four backdoor attack types demonstrate CLIP-Fed’s superiority. Compared to state-of-the-art methods, CLIP-Fed achieves an average reduction in ASR, i.e., 2.03% on CIFAR-10 and 1.35% on CIFAR-10-LT, while improving average MA by 7.92% and 0.48%, respectively. Additionally, CLIP-Fed maintains ̵client-side efficiency without extra computational overhead beyond local training.
Related Work
Backdoor Defense in FL. To mitigate the backdoor attacks, for example, Auror (Shen, Tople, and Saxena 2016), Flame (Nguyen et al. 2022), FedDMC (Mu et al. 2024), and FreqFed (Fereidooni et al. 2024) tried model clustering to eliminate malicious backdoors while ensuring the performance of the aggregated model. The effectiveness of such methods depended on an assumption of homogeneous client data distribution, such that the performance under Non-IID conditions is limited. Other methods, e.g., BRCA (Zhai et al. 2021) and FLTrust (Cao et al. 2021), assign trust scores to clients using a small clean validation set hosted on the server to detect and filter malicious contributions. While effective in some cases, this paradigm contradicts the decentralized and privacy-preserving nature of FL by requiring centralized access to client data. Recent work has explored contrastive learning as a defense mechanism. Through contrastive learning (Radford et al. 2021), FLPurifier (Zhang et al. 2024) severs latent correlations between target labels and trigger patterns, thereby neutralizing potential backdoor artifacts during model training phases. Although this strategy improves model robustness, it increases the computational burden on the client and makes it less suitable for resource-constrained environments. Overall, achieving effective backdoor defense under Non-IID conditions while ensuring privacy and avoiding increasing client-side overhead remains a critical challenge. Our scheme proposes a server-side dataset augmentation strategy based on MLLMs (Lee et al. 2024) and frequency analysis, which enhances the dataset’s coverage against diverse backdoor triggers without accessing client data. Additionally, our scheme integrates a global purification mechanism combining prototype contrastive learning and knowledge distillation, such that effective suppression of backdoor patterns in Non-IID FL can be achieved.
Figure 2: The framework of the proposed CLIP-Fed. CLIP-Fed contains four modules: Data Augmentation for Server Dataset aims to build the server dataset, Dynamic Model Filtering by Clustering aims to filter out malicious models before aggregation, Feature Rectification via Prototype Alignment and Global Model Knowledge Transfer aim to purify the backdoor.
Vision-Language Pre-train Model. VLPs (Radford et al. 2021; Li et al. 2022; Liu et al. 2023) have made significant strides across various downstream tasks such as image classification, visual question answering, and cross-modal retrieval. CLIP (Radford et al. 2021) is a representative VLP that uses contrastive learning to capture semantic relationships between images and texts, enabling both to be mapped into a shared feature space. Recent studies have also explored integrating contrastive learning into FL to address the challenges of Non-IID client data. MOON (Li, He, and Song 2021) proposes a model-contrastive FL framework that utilizes inter-client representation similarity to improve model generalization under heterogeneous data distributions. In addition, the strong semantic understanding of VLPs has shown promise in handling long-tailed distributions in FL (Shi et al. 2024), as VLPs can capture rich visual concepts even from underrepresented classes. However, while such capabilities indicate that VLPs are well suited for learning robust and semantically meaningful representations, their potential for enhancing security in FL, particularly in defending against backdoor attacks, remains underexplored. To the best of our knowledge, our work is the first to use the VLPs for backdoor defense in FL. Leveraging the cross-modal representation capacity of VLPs, we align the global model’s visual features with those of a VLP to detect and eliminate backdoor patterns, establishing a novel and practical defense paradigm.
Methodology
Problem Formulation
In the FL setup with one server and $m$ clients, each client $i$ holds a local dataset $D_{i}$ and a model $w_{i}=(f_{i},c_{i})$, where $f_{i}$ is the feature extractor and $c_{i}$ is the classifier. The server maintains a global model $G_{s}=(f_{s},c_{s})$. For an input $x$, the prediction is obtained via $p=c(f(x))$. Malicious clients conduct backdoor attacks by embedding triggers in training samples and uploading poisoned models. The manipulated models aim to cause the global model to misclassify triggered inputs as the target label $y^{\prime}$. The Attack Success Rate (ASR) is defined as:
$\mathrm{ASR}=\underset{\begin{subarray}{c}(x,y)\sim D\\ y\neq y^{\prime}\end{subarray}}{\mathbb{E}}\left[\operatorname{Pr}\left(\Phi(t(x))=y^{\prime}\right)\right],$ (1)
where $D$ is the dataset for the FL task, $y$ is the label for input $x$, $y^{\prime}$ is the target label, and $t(\cdot)$ is the backdoor trigger function. The framework of CLIP-Fed is shown in Figure 2. CLIP-Fed aims to suppress backdoor and reduce ASR.
Data Augmentation for Server Dataset
Generation of Image-Text Samples
In CLIP-Fed, to preserve client data privacy, the server cannot directly utilize client datasets for contrastive learning and knowledge distillation. MLLMs can be utilized to perform server-side dataset generation and augmentation without accessing client data (Lee et al. 2024; Pan et al. 2025). Unlike traditional methods that rely on geometric transformations or noise injection, our scheme semantically expands candidate labels to guide the image generation. For instance, combining object descriptors with contextual modifiers such as “a small airplane beneath a cloudy sky” or “a frog partially obscured by leaves” naturally introduces variations in scale, occlusion, and background complexity, which inherently influence visual perception. The synthesized images are then adaptively resized and normalized to match the statistical distribution of the target dataset, such that the alignment with the original feature space is guaranteed.
Frequency Sensitivity-guided Image Augmentation
To enhance the robustness of backdoor purification to triggers, we add perturbations to image samples of the server dataset to simulate images with triggers while preserving label information. Specifically, we first add visible triggers to a clean image $x$ to obtain $\bar{x}$. Then, we apply the Discrete Cosine Transform (DCT) to both $x$ and $\bar{x}$ and divide the spectra into multiple frequency bands. We then identify the frequency intervals most sensitive to triggers by comparing the similarities between images $x$ and $\bar{x}$ in different frequency domains. We assume that frequency intervals with the lowest similarity have the highest sensitivity to triggers. The DCT spectrogram shows low-frequency coefficients clustered in the upper-left and high-frequency components in the lower-right, symmetric about the main diagonal. We segment the spectrogram into discrete frequency bands ordered from low to high, aligned orthogonally to the diagonal. Specifically, the matrix is unevenly divided into $K$ segments along the diagonal, with the distance from the $k$-th division point to the top-left corner set to $1/2^{(K-k)}$ of the diagonal length. We then extend these $K$ frequency bands into $K$ independent spectrograms, represented as $\left(F_{1}(u,v),\ldots,F_{K}(u,v)\right)$. We assess the distance between the spectrogram coefficients in each frequency segment using Mean Squared Error (MSE). Frequency regions with higher MSE values exhibit lower similarity, indicating higher sensitivity to pixel-block triggers. The frequency band with the highest MSE is denoted as $F_{high}(u,v)$, where we inject Gaussian noise perturbations into the original image. The image is then reconstructed into the spatial domain via IDCT, yielding an enhanced server dataset image with noise added only in the sensitive $F_{high}$ region. Perturbing only the most sensitive frequency bands better simulates backdoor triggers while minimizing impact on model utility.
Dynamic Model Filtering by Clustering
Clustering presents a potential solution for distinguishing between malicious and benign models in FL. In CLIP-Fed, we adopt HDBSCAN for pre-aggregation filtering, following advanced methods (Nguyen et al. 2022). Because high-dimensional parameter spaces suffer from the curse of dimensionality, which reduces the reliability of distance metrics like cosine similarity, we use PCA to project model parameters into a lower-dimensional space where distance computations are more meaningful and robust. Following the threat assumptions and parameter settings established in Flame (Nguyen et al. 2022) and FreqFed (Fereidooni et al. 2024), we operate under the premise that benign models constitute the majority of clients while malicious models remain in the minority. Accordingly, we set the minimum cluster size to encompass at least 50% of clients, ensuring the resulting clusters contain predominantly benign model updates. While this clustering process enables CLIP-Fed to eliminate most malicious models prior to aggregation, it is acknowledged that the identified benign clusters may still contain some residual malicious models. After model clustering, CLIP-Fed performs global model aggregation as specified in:
$\theta_{s}^{t+1}=\sum_{i\in S_{t}}\frac{D_{i}}{D_{S_{t}}}\theta_{i}^{t},\ \theta\in\{f,c\},$ (2)
where $S_{t}$ denotes the set of clients in the clustering cluster. In each training round, CLIP-Fed dynamically selects different clients to participate in aggregation, thereby resisting attacks from adaptive adversaries who randomly compromise clients during each round. The framework subsequently employs post-aggregation backdoor elimination strategies to purify further any remaining backdoor patterns that may persist in the global model.
Datasets Attack FedAvg FLTrust FLAME FoolsGold FEDCPA CLIP-Fed
MA/ASR(%) MA/ASR(%) MA/ASR(%) MA/ASR(%) MA/ASR(%) MA/ASR(%)
CIFAR-10 Badnets 78.87 / 96.11 73.53 / 84.78 70.75 / 3.55 72.66 / 83.87 78.65 / 38.37 81.50 / 2.77
DBA 78.23 / 91.69 76.24 / 46.69 72.75 / 4.22 77.83 / 56.29 77.57 / 35.05 81.75 / 1.86
LayerAttack 79.30 / 14.18 76.30 / 1.42 72.50 / 2.54 67.39 / 8.38 78.39 / 1.77 76.70 / 0.19
AFA 77.43 / 18.05 76.42 / 6.96 73.18 / 4.05 70.77 / 14.03 76.23 / 4.00 80.91 / 1.42
CIFAR-10-LT Badnets 70.88 / 80.17 63.83 / 64.05 69.13 / 5.16 66.46 / 40.27 71.96 / 4.16 73.07 / 4.01
DBA 70.83 / 62.00 62.22 / 40.22 69.21 / 4.62 65.24 / 66.02 72.36 / 3.47 72.73 / 3.38
LayerAttack 72.30 / 14.18 68.06 / 8.73 68.69 / 7.58 66.15 / 15.38 71.17 / 3.82 70.00 / 0.29
AFA 71.67 / 17.18 67.45 / 5.19 67.34 / 4.34 47.95 / 7.29 72.43 / 3.86 74.05 / 2.25
CIFAR-100 Badnets 43.06 / 97.12 36.07 / 48.48 45.09 / 38.02 37.45 / 88.75 47.68 / 1.19 45.47 / 0.39
DBA 45.96 / 93.68 41.38 / 56.88 45.24 / 0.64 44.90 / 95.15 48.29 / 1.19 46.17 / 0.39
LayerAttack 46.40 / 10.48 43.63 / 0.98 41.59 / 1.74 47.68 / 2.40 49.48 / 0.52 44.94 / 0.00
AFA 41.82 / 13.51 42.64 / 0.51 44.91 / 0.62 44.01 / 2.64 47.61 / 0.59 44.57 / 0.30
Table 1: Performance comparison between CLIP-Fed and other defense methods.
Feature Rectification via Prototype Alignment
In this phase, CLIP-Fed utilizes the CLIP model, which is equipped with rich prior knowledge, as an auxiliary module to guide the rectification and retraining of the feature extraction module of global model on the server. In backdoor models, the feature extraction module exhibits the increased sensitivity to trigger-associated patterns. Existing research (Shi et al. 2024) has shown that leveraging CLIP’s strongly supervised feature prototypes can effectively guide the generation of federated features. Inspired by this insight, we employ contrastive learning between CLIP-derived class prototypes and features generated by the global model feature extractor, thus mitigating prototype deviations caused by backdoor trigger interference. As shown in Eq. 3, in the first part, we treat the output of the feature extractor $v_{i}^{c}$ (Sample $i$ belonging to class $c$) and the output of CLIP’s visual encoder $e_{c,\text{clip}}^{I}$ as positive pairs, while considering the output of the feature extractor $v_{i}^{c}$ and other class prototypes $e_{k,\text{extractor}}$ from the feature extractor as negative pairs.
$\mathcal{L}_{i}^{\text{proto-1}}=-\log\frac{\exp\left(sim\langle v_{i}^{c},e_{c,\text{clip}}^{I}\rangle/\tau\right)}{\sum_{k=1[k\neq c]}^{C}\exp\left(sim\langle v_{i}^{c},e_{k,\text{extractor}}\rangle/\tau\right)},$ (3)
where $sim\left\langle\cdot\right\rangle$ denotes cosine similarity and $\tau$ is a temperature hyperparameter. As shown in Eq. (4), in the second part, we form positive pairs between the feature extractor’s output $v_{i}^{c}$ and outputs of the CLIP’s text encoder $e_{c,\text{clip}}^{T}$, with negative pairs being the feature extractor’s output $v_{i}^{c}$ and other class prototypes $e_{k,\text{clip}}^{T}$ from the CLIP’s text encoder.
$\mathcal{L}_{i}^{\text{proto-2}}=-\log\frac{\exp\left(sim\langle v_{i}^{c},e_{c,\text{clip}}^{T}\rangle/\tau\right)}{\sum_{k=1[k\neq c]}^{C}\exp\left(sim\langle v_{i}^{c},e_{k,\text{clip}}^{T}\rangle/\tau\right)}.$ (4)
The total prototype contrastive loss as shown in:
$\mathcal{L}_{\text{pcl}}=\sum_{i=1}^{C\times m}\left(\mathcal{L}_{i}^{\text{proto-1}}+\mathcal{L}_{i}^{\text{proto-2}}\right).$ (5)
The proposed approach systematically eliminates backdoor-induced deviations in feature representations while maintaining the model’s core classification performance.
Global Model Knowledge Transfer
To thoroughly eliminate residual backdoor patterns in the global model, CLIP-Fed employs knowledge transfer to purify the global model using CLIP’s prior knowledge. In this framework, CLIP acts as a “teacher” model, while the global model serves as the “student” model. During each iteration, CLIP provides guidance to the global model using the server dataset, progressively reducing the impact of the backdoor. We utilize the KL divergence to measure the discrepancy between logits distributions of the global model and CLIP on the annotated dataset. Specifically, the loss function for the knowledge transfer process is formulated as
$\mathcal{L}_{\text{kt}}=\mathcal{L}_{ce}\left(y,l_{\text{global}}\right)+\beta\cdot\mathcal{D}_{\text{KL}}\left(l_{\text{clip}}\|l_{\text{global}}\right),$ (6)
where $l_{\text{global}}$ and $l_{\text{clip}}$ denote the output logits of the global model and CLIP. This loss of distillation ensures that the global model aligns its predictions with CLIP’s robust, backdoor-free representations, thus eradicating malicious patterns while preserving model utility.
Experiments
Datasets Attack w/o FR w/o KD CLIP-Fed
MA/ASR MA/ASR MA/ASR
CIFAR-10 BadNets 75.88/6.30 75.08/7.37 81.43/2.62
DBA 76.57/3.31 75.73/4.01 81.56/2.05
LayerAttack 75.43/6.15 73.52/5.39 77.10/0.19
AFA 76.17/3.52 73.79/4.53 80.94/1.30
CIFAR-10-LT BadNets 70.88/8.87 71.31/6.93 73.40/4.11
DBA 70.32/4.30 70.76/5.66 73.00/3.29
LayerAttack 70.10/4.20 69.55/2.15 70.20/0.38
AFA 70.87/4.48 71.55/3.41 73.86/2.16
CIFAR-100 BadNets 44.68/0.63 44.83/0.45 45.56/0.41
DBA 44.35/0.63 45.44/0.45 45.79/0.31
LayerAttack 43.97/0.19 41.21/0.00 44.51/0.00
AFA 45.50/0.57 44.99/0.34 45.61/0.22
Table 2: Performance comparison between CLIP-Fed w/o FR, CLIP-Fed w/o KD, and CLIP-Fed.
(a) CLIP
(b) BadNets without defense
(c) BadNets with CLIP-Fed 50 epochs
(d) BadNets with CLIP-Fed 200 epochs
Figure 3: The visualization of the feature representations under different attacks on CIFAR-10-LT with/without CLIP-Fed.
(a) MA with 20% malicious clients
(b) MA with 30% malicious clients
(c) MA with 40% malicious clients
(d) ASR with 20% malicious clients
(e) ASR with 30% malicious clients
(f) ASR with 40% malicious clients
Figure 4: Comparison of MA and ASR against BadNets attacks with varying malicious client proportions under CIFAR-10-LT.
Proportions of
Malicious Clients
FedAvg FLTrust FLAME FoolsGold FEDCPA CLIP-Fed
MA/ASR(%) MA/ASR(%) MA/ASR(%) MA/ASR(%) MA/ASR(%) MA/ASR(%)
20% 71.21 / 89.74 64.17 / 61.93 69.13 / 5.16 67.85 / 43.99 70.17 / 5.52 73.74 / 4.54
30% 70.62 / 89.81 62.39 / 67.40 67.40 / 8.17 68.25 / 74.94 69.61 / 90.22 72.37 / 4.70
40% 67.98 / 97.63 66.40 / 85.97 66.49 / 10.59 66.71 / 78.41 67.01 / 96.54 72.15 / 4.93
Table 3: Performance comparison against BadNets attacks with different proportions of malicious clients under CIFAR-10-LT.
Degree of
Non-iid
CIFAR-10 CIFAR-100
No Defense CLIP-Fed No Defense CLIP-Fed
MA/ASR MA/ASR MA/ASR MA/ASR
0.2 78.51/97.04 82.41/2.21 45.06/97.12 45.72/0.77
0.4 77.25/95.63 81.50/2.77 44.00/94.49 45.68/0.44
0.6 77.42/95.86 78.23/0.90 43.57/91.05 45.60/0.32
Table 4: Performance of CLIP-Fed against BadNets attacks with different degrees of Non-IID.
Poisoning Ratio 60% 80% 100%
Defense No Defense CLIP-Fed No Defense CLIP-Fed No Defense CLIP-Fed
Trigger Size MA/ASR(%) MA/ASR(%) MA/ASR(%) MA/ASR(%) MA/ASR(%) MA/ASR(%)
$\times1$ 70.06 / 11.68 71.83 / 3.92 67.34 / 67.71 71.33 / 4.67 71.21 / 89.74 72.49 / 4.16
$\times2$ 70.45 / 15.43 72.12 / 4.86 69.15 / 77.74 71.69 / 4.07 71.05 / 94.37 72.11 / 5.40
$\times3$ 69.56 / 20.08 70.93 / 2.68 68.50 / 88.77 72.58 / 2.34 70.60 / 95.55 72.57 / 2.16
$\times4$ 69.51 / 28.51 72.53 / 3.99 69.42 / 89.72 72.74 / 2.55 70.40 / 96.11 72.95 / 3.67
Table 5: Defense performance against BadNets attacks with different attack strength under CIFAR-10-LT.
Datasets. We evaluated CLIP-Fed on CIFAR-10, CIFAR-10-LT, and CIFAR-100 (Krizhevsky 2009). CIFAR-10-LT introduced an exponential class imbalance to simulate real-world long-tailed distributions. To emulate Non-IID client settings, we adopted Pathological Non-IID partitioning (Blanchard et al. 2017).
Implementation Details. We simulated a FL scenario with 100 clients, where a portion is malicious. Malicious clients conducted four types of backdoor attacks, including BadNets (Gu, Dolan-Gavitt, and Garg 2017), DBA (Xie et al. 2020), LayerAttack, and AFA, following an “all-to-one” strategy. ResNet-18 was used as the model backbone. All experiments were conducted on an RTX 4090 GPU with Python 3.9, PyTorch 2.2.2, and CUDA 11.8. Unless stated otherwise, we simulated Non-IID client data distributions and employed the ViT-L/14 version of CLIP.
Baselines and Existing Defenses. We compared CLIP-Fed with representative backdoor defense methods in FL, including: (1) FedAvg (McMahan et al. 2017), a standard aggregation baseline without defense, (2) FLAME (Nguyen et al. 2022) that combines clustering and differential privacy, (3) FLTrust (Cao et al. 2021) that is a server-side validation-based defense, (4) FoolsGold (Fung, Yoon, and Beschastnikh 2020), a defense scheme based on the similarity of contributions among clients, and (5) FEDCPA (Han et al. 2023) that focuses on critical parameter similarity.
Performance Comparison
Table 1 presented a comparison of Main Task Accuracy (MA) and Attack Success Rate (ASR) across different methods under varying data distributions and backdoor attack. The results demonstrated that CLIP-Fed consistently achieved effective mitigation of backdoor across various attacks and datasets. Specifically, CLIP-Fed achieved the lowest ASR under all four attack types on both the CIFAR-10 and CIFAR-10-LT datasets. On the CIFAR-10, our scheme achieved ASR reductions of 0.78%, 2.36%, 2.35%, and 2.63% under BadNets, DBA, LayerAttack, and AFA, respectively, compared to the SOTA defense FLAME. On the CIFAR-10-LT dataset that exhibited a significant long-tail distribution, our scheme reduced ASR by 0.15%, 0.09%, 3.53%, and 1.61% under the same four attacks, compared to the SOTA defense FEDCPA. In terms of MA, our scheme also had advantages over existing solutions, as it achieved backdoor defense without significantly compromising task performance. This indicated that CLIP-Fed was able to accurately filter out malicious models while preserving the contributions of benign ones, making it more effective than standalone pre-aggregation detection or post-aggregation noise injection methods. On the CIFAR100 dataset with a higher-level complexity of classification tasks, CLIP-Fed achieved the best defense under the four types of attacks.
Ablation Study
We evaluated two variants of CLIP-Fed: CLIP-Fed w/o FR (without feature extractor rectification), and CLIP-Fed w/o KD (without knowledge distillation). Table 2 compared the accuracy of the global model and the defense performance under identical experiment setups. We observed that both CLIP-Fed w/o FR and CLIP-Fed w/o KD achieved lower ASR than FedAvg (as shown in Table 1) across all three datasets and four types of attacks. The results indicated that individually applying either feature extractor retraining or global model knowledge distillation could partially reduce the association between trigger patterns and target labels, thereby mitigating backdoor effects. Moreover, CLIP-Fed combined both feature retraining and knowledge distillation, such that it achieved the highest MA and lowest ASR across all datasets and attack methods.
Visualization of Backdoor Elimination
As shown in Figure 3, t-SNE was used to visualize the feature space distribution of test samples under BadNets attacks, both with and without CLIP-Fed. We added triggers to 5% of the test samples. In the visualization, samples containing triggers were outlined in black. Figure 3a presented the visualization of test samples after being processed by the CLIP model. The CLIP model demonstrated strong robustness to trigger samples and successfully clustered samples from the same class together in the feature space. This suggested that using CLIP’s prior knowledge to guide FL in backdoor purification was a feasible approach. Figure 3b showed the feature space clustering after a backdoor attack without any defense. We observed that nearly all samples with triggers formed a separate cluster, indicating a successful backdoor attack. Figures 3c and 3d showed the clustering results after training with CLIP-Fed for 50 and 200 epochs, respectively. The trigger samples were no longer isolated and rejoined the clusters of clean samples from the same class. The results indicated that CLIP-Fed disrupted backdoor feature clustering and weakened the link between triggers and target labels, effectively mitigating the backdoor.
Parameter Impact
The Impact of Number of Malicious Clients. Table 3 and Figure 4 presented a comparison of the defense performance of CLIP-Fed and existing methods under varying proportions of malicious clients. When the proportion of malicious clients was 0.2, FEDCPA, FLAME, and CLIP-Fed demonstrated comparable effectiveness in reducing the ASR. However, as the proportion exceeded 0.2, the defense performance of existing methods began to degrade to varying extents. Notably, FEDCPA experienced the most significant drop in performance, with an ASR of 90.22% when the malicious client ratio was 0.3, and 96.54% when it increased to 0.4. In contrast, CLIP-Fed consistently maintained strong defense performance across all settings. When the proportion of malicious clients was 0.3 and 0.4, the ASR of CLIP-Fed remained significantly lower than that of the SOTA method FLAME by 3.47% and 5.66%, respectively. In terms of MA, CLIP-Fed also performed favorably, surpassing FedAvg by 2.53%, 1.75%, and 4.17% at malicious client ratios of 0.2, 0.3, and 0.4, respectively. A single pre-aggregation defense strategy made existing methods vulnerable to high proportions of malicious clients, whereas CLIP-Fed’s post-aggregation backdoor purification supplements the defense capability under such conditions. Overall, compared with existing approaches, CLIP-Fed defended against backdoor attacks across different malicious client ratios while minimizing the degradation of task performance.
The Impact of Non-IID. Table 4 presented the defense performance of CLIP-Fed against BadNets attack under varying degrees of Non-IID data distribution. On the CIFAR-10 dataset, the CLIP-Fed approach achieved an average MA of 80.71% and an average ASR of 1.96%. On the CIFAR-100 dataset, the average MA was 45.67% and the average ASR was 0.51%. Across both datasets, the degree of Non-IID distribution caused only minor fluctuations in MA and ASR, and its impact on defense performance was negligible. This was because heterogeneous client data distributions primarily affected the filtering of malicious models before aggregation, while CLIP-Fed’s backdoor purification strategy was decoupled from model aggregation, thereby eliminating the impact of Non-IID data on purification effectiveness. The above demonstrated that CLIP-Fed was effective under different levels of Non-IID data distribution, maintaining high MA and low ASR for robust backdoor defense.
Defense under Different Attack Strength. Trigger size and poisoning ratio served as crucial factors affecting the attack strength. As shown in Table 5, in the absence of defense, MA gradually declined while ASR increased as the trigger size became larger. Larger triggers resulted in more distinguishable features, making it easier for the model to associate them with the target label. Additionally, higher poisoning rates further intensified the attack, leading to a sharper drop in model performance. When CLIP-Fed was applied, backdoor were effectively prevented from being embedded into the global model, regardless of trigger size or poisoning rate. The connection between trigger features and the target label was severed, and residual backdoor patterns in the global model were removed, making the defense independent of attack strength. The above findings indicated that the influence of trigger size and poisoning rate became minimal when CLIP-Fed was employed.
Conclusion
In this paper, we proposed CLIP-Fed that is a general backdoor defense scheme guided by the VLP model CLIP. Without accessing client data or assuming homogeneous distributions, CLIP-Fed enhanced effectiveness by combining pre-aggregation malicious model filtering with post-aggregation purification. Leveraging MLLMs and frequency analysis, we constructed and enhanced a server dataset, addressing privacy concerns and improving coverage of diverse backdoor triggers with minimal utility loss. We further proposed a global model purification strategy based on contrastive learning and knowledge distillation, which corrected class prototype deviations and weakened the link between triggers and target labels, thus improving robustness against trigger samples. Extensive experiments on multiple datasets demonstrated that CLIP-Fed outperformed existing methods in defending against backdoor attacks under varying data distributions and malicious proportions of clients.